OpenStudy (anonymous):

3(6-4x)=2(-6x+9) ste by step plz

OpenStudy (agreene):

start by distributing 18-12x=-12x+18 we can rewrite 18-12x=18-12x <== this is a true statement x= any real number infinite solutions.

OpenStudy (agreene):

you can take that further, to 0=0 if you choose.

OpenStudy (anonymous):

3(6-4x) = 2(-6x + 9) Always remember to translate between Mathematics & what you understand One side must be equal to the other side 1) Distribute 18 -12x = -12x + 18 2) Combine Like terms 18 -18 = -12x +12x 3) Do the operations asked 0 = 0 *Translate Zero is equal to Zero therefore the statement is TRUE GOOD LUCK!